OpenStudy (anonymous):

Write 2/5/7 as an improper fraction.

OpenStudy (anonymous):

37/7

OpenStudy (anonymous):

do you know how i got 37/7?

OpenStudy (anonymous):

Magic?

OpenStudy (anonymous):

how? i'm so bad at math

OpenStudy (anonymous):

\[\frac{ \frac{ 2 }{ 5 } }{ 7 } = \frac{ 2 }{ 5 }*\frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

Then multiply across the top and the bottom.

OpenStudy (anonymous):